OpenStudy (anonymous):

What is 15/25 as a percent

OpenStudy (anonymous):

percent means out of 100 what do you have to multiply 25 by to get 100. multiply 15 by that same amount and that is the percent. post what you get and i'll check it.

OpenStudy (anonymous):

Isn't the answer 60 percent

OpenStudy (anonymous):

With what you told me about how to do 15/25 do u do The same for 17/25

OpenStudy (anonymous):

15*4/25*4 = 60/100 = 60% yes, you do the same thing with 17/25

OpenStudy (anonymous):

Thank you so much math is very hard for me

OpenStudy (anonymous):

your welcome. if you are having trouble with math you might want to check out khanacademy.org they have very good tutorials